# LitterBox+: An Extensible Framework for LLM-enhanced Scratch Static Code Analysis

### v1

@@ Abstract
Index Terms:
Scratch, Block-based Programming, LLM, Automated Feedback
@@ II Background
The different shapes of Scratch blocks determine how they can be combined into programs, thus simplifying programming. However, learners may still struggle to correctly apply programming concepts and produce buggy code. While the process of finding and fixing bugs is an inherent part of learning to program, automated tools can offer support. Since Scratch code is block-based, dedicated analysis tools are required. For example, LitterBox [4] is a linter that automatically detects bugs based on recurring patterns that are common manifestations of misconceptions. Figure 1 shows such a bug pattern in the popular ‘BoatRace’ program [^1] https://scratch.mit.edu/projects/63957956/, accessed 2025-07-21 where a misconception resulted in a ‘Missing Loop’ bug pattern, resulting in a collision detection only being executed once at the start of program execution. Such bug patterns are common in Scratch [3], and feedback on them has been shown to support learners [5]. However, classic static analysis tools like LitterBox can only provide generic help, and learners have to rely on other sources for clarifications or suggestions on how to fix their code.
@@
III System Architecture
Figure 2: LitterBox+ architecture overview.
The main interface to our tool for students is via an extension to the Scratch user interface. This direct integration ensures a low barrier to entry by not requiring context switches away from the familiar programming environment. We added a toggleable panel that shows by default issues found by the regular LitterBox static code analysis in the current program (cf. Fig. 1). For each issue, the user can request an extended llm-generated explanation, or request a fix for the issue, which will be automatically integrated into the active program. Additionally, our gui extension offers a chat-like input where a user can ask arbitrary questions about the current sprite or program.
To support the communication between LitterBox running as a Java application on a server and the website in the browser, we extend LitterBox with a rest-api. When we receive a request relating to an issue found by LitterBox from the user interface, we combine information from the issue itself and the program to construct a prompt for the llm. From the issue itself, we extract its type (e.g., Missing Loop) and the default description added by LitterBox. To embed the issue in the relevant context, we also extract the source code of the Scratch program. To provide this code to the llm, we convert the program into the textual scratchblocks format. We use the same format as is used for example on the Scratch community forums and only extend it slightly to include special comments adding unique ids to the start of sprites and scripts. The scratchblocks format was likely seen by the llm during training due to its use on the forums, thus resulting in improved performance over the json program representation. Unlike taking screenshots of the workspace, this removes the visual component from the code. This avoids the problem that code blocks are allowed to overlap in the Scratch coding area; by using the textual format instead of a screenshot of the code, we therefore have more control over which part of the code is included in our prompt.
IV User Interfaces
The main user interface is integrated directly into the Scratch web interface. While tests during the implementation of our tool showed that llms can understand and reason about Scratch code, the model may still occasionally produce misleading or wrong explanations and fixes. Since especially learners often cannot discern such cases from helpful ones, we show a prominent hint about the potentially wrong nature in every location where llm-generated content is presented. For example, the red exclamation mark in Fig. 1 shows this hint as tooltip when hovered over. Even though our tool is llm-agnostic, we call all llms GPT in the user interface, since students are likely more familiar with GPT as a general term for all kinds of llms.
@@ IV User Interfaces
Finally, to allow future research to extend upon our framework, all LitterBox functionality is available via its Java api so that it can also be integrated as a library into other tools. This api allows for future extension in various directions: (1) Our prompt designs can be replaced by customised prompt templates. (2) Future extensions could add support for more model providers besides the currently supported externally hosted (OpenAI) and self-hosted (Ollama) models. [^3] Our implementation uses langchain4j, https://docs.langchain4j.dev/, accessed 2025-07-14. (3) New llm-based features that make use of the existing static code analysers and the scratchblocks parser, or ones based on entirely new analysers could be introduced.
@@ V-B Explanation for LitterBox Issues
[…]
@@ V-C Automatic Issue Fixing
To extend the support beyond providing an more specific description of the issue, our framework can also propose fixed versions of the code and integrate them directly into the current program. Similar to the explanation, we added a ‘GPT: Fix!’ button into the user interface next to the shown LitterBox issue.
@@
[…]
To integrate the parsed scripts into existing programs, we rely on the special comments marking the beginning of sprites and scripts (cf. Section III). Using the sprite names and block ids contained in these comments, we can align the parsed code snippets with the original code and make the required replacements. These replacements are performed on the internal program representation of LitterBox (i.e., the abstract syntax tree), which is substantially more flexible and reliable compared to working directly on the textual scratchblocks- or json-formats. Any sprites or scripts with unseen ids are added as new code to the program. When adding a new sprite, we use the default image of the Scratch cat mascot.The updated program is finally converted by LitterBox back into the Scratch-native json format and loaded into the Scratch-gui where the user can inspect it in the native code editor (upper half of Fig. 1) and continue to edit it. In case the proposed fix is not satisfactory, the user can choose to revert the change.
@@ VI Future Work
Since we introduce an initial version of our extensible LitterBox+ framework in this paper, this leaves various opportunities for future work. For use in a classroom setting where the target solution of the program is known, additional agentic capabilities could be integrated that automatically run Whisker [9] or block-based [10] test suites so that either llm-proposed code analysis issue fixes can be checked and improved automatically via a follow-up prompt to the llm, or that fixes for failing tests can be requested directly by the user. Since the Scratch-gui is lacking any kind of code completion feature at the moment, our llm-based completion could be a valuable further extension to the framework and its users. Finally, all features proposed as part of LitterBox+ should be evaluated with children to understand further in how such llm-based tools are used [6, 7] to then adjust prompts and user interface, and develop new features based on the findings.

### v2

@@ II Background
The different shapes of Scratch blocks determine how they can be combined into programs, thus simplifying programming. However, learners may still struggle to correctly apply programming concepts and produce buggy code. While the process of finding and fixing bugs is an inherent part of learning to program, automated tools can offer support. Since Scratch code is block-based, dedicated analysis tools are required. For example, LitterBox [4] is a linter that automatically detects bugs based on recurring patterns that are common manifestations of misconceptions. Figure 1 shows such a bug pattern in the popular ‘BoatRace’ program [^1] https://scratch.mit.edu/projects/63957956/, accessed 2025-07-21 where a misconception caused a ‘Missing Loop’ bug pattern, resulting in a collision detection only being executed once at the start of program execution. Such bug patterns are common in Scratch [3], and feedback on them has been shown to support learners [5]. However, classic static analysis tools like LitterBox can only provide generic help, and learners have to rely on other sources for clarifications or suggestions on how to fix their code.
@@
III User Interfaces
The main user interface is integrated directly into the Scratch web interface. This direct integration ensures a low barrier to entry by not requiring context switches away from the familiar programming environment. We added a toggleable panel that shows by default issues found by the regular LitterBox static code analysis in the current program (cf. Fig. 1). For each issue, the user can request an extended llm-generated explanation, or request a fix for the issue, which will be automatically integrated into the active program. Additionally, our gui extension offers a chat-like input where a user can ask arbitrary questions about the current sprite or program.
While early trials of our tool showed that llms can understand and reason about Scratch code, the model may still occasionally produce misleading or wrong explanations and fixes. Since especially learners often cannot discern such cases from helpful ones, we show a prominent hint about the potentially wrong nature in every location where llm-generated content is presented. For example, the red exclamation mark in Fig. 1 shows this hint as tooltip when hovered over. Even though our tool is llm-agnostic, we call all llms GPT in the user interface, since students are likely more familiar with GPT as a general term for all kinds of llms.
@@ III User Interfaces
Finally, to allow future research to extend upon our framework, all LitterBox functionality is available via its Java api so that it can also be integrated as a library into other tools. This api allows for future extension in various directions: (1) Our prompt designs can be replaced by customised prompt templates. (2) While the user can already configure LitterBox+ to use either externally hosted (OpenAI) or self-hosted (Ollama) models, additional model providers could be added. [^3] Our implementation uses langchain4j, https://docs.langchain4j.dev/, accessed 2025-07-14. (3) New llm-based features that make use of the existing static code analysers and the scratchblocks parser, or ones based on entirely new analysers could be introduced.
@@
IV System Architecture
Figure 2: LitterBox+ architecture overview.
To support the communication between LitterBox running as a Java application on a server and the website in the browser, we extend LitterBox with a rest-api. When we receive a request relating to an issue found by LitterBox from the user interface, we combine information from the issue itself and the program to construct a prompt for the llm. From the issue itself, we extract its type (e.g., Missing Loop) and the default description added by LitterBox. To embed the issue in the relevant context, we also extract the source code of the Scratch program. To provide this code to the llm, we convert the program into the textual scratchblocks format. We use the same format as is used for example on the Scratch community forums and only extend it slightly to include special comments adding unique ids to the start of sprites and scripts. The scratchblocks format was likely seen by the llm during training as it is common on forums and uses a sequential structure for code, unlike Scratch’s internal json program representation, thus resulting in improved performance. Unlike taking screenshots of the workspace, this removes the visual component from the code. This avoids the problem that code blocks are allowed to overlap in the Scratch coding area; by using the textual format instead of a screenshot of the code, we therefore have more control over which part of the code is included in our prompt.
@@ V-B Explanation for LitterBox Issues
1. How the faulty code affects observable behavior of sprites:The static code analysis tool pointed out that a certain condition is onlychecked once, instead of being checked continuously.(a) This script will only check once, immediately after the green flag isclicked, if the sprite is touching the color .(b) If at that exact moment the sprite is NOT touching this color,nothing will happen, even if the sprite touches it later.(c) As a result, the response (move -1 steps) won’t trigger when thesprite later comes into contact with the specified color, making thecode seem unresponsive.(d) Observable Effect: The Boat sprite will NOT react (move -1 steps)when it touches the color at any point after the game starts,unless it happens to be touching it exactly as the green flag is clicked.2. List of steps to observe the faulty behavior:1. Press the green flag to start the program.2. Move the Boat sprite (using code or manually, depending on yoursetup) so that it touches the color after the game has started.3. Observe: The Boat sprite does not move backwards (-1 steps),despite touching the color.4. (Optional) Restart and have the sprite start on top of color. Now,you might see it move -1 steps at game start, since the condition wastrue at that moment.
@@ V-C Automatic Issue Fixing
To extend the support beyond providing a more specific description of the issue, our framework can also propose fixed versions of the code and integrate them directly into the current program. Similar to the explanation, we added a ‘GPT: Fix!’ button into the user interface next to the shown LitterBox issue.
@@
While we observed that the llms can in most cases produce valid scratchblocks code, they still often repeat similar issues, like for example using a non-existing block set rotation to Àg instead of the actual point in direction Àg or using braces to scope inner blocks of control statements instead of using the correct then and end markers. Sometimes the llm also includes both the buggy and the fixed scripts in its output with the script marker comments (cf. Section IV) modified with an additional ‘original version’ or ‘modified version’ suffix, even though our prompts explicitly request the llm to not modify the special comments. To mitigate such frequent problems, we introduced a postprocessing step to the llm’s output that applies heuristic rules fixing common syntactical issues before passing the scratchblocks code to the actual parser. In case this parsing still fails, we prompt the llm with only the unparseable scripts again in up to three attempts to request a syntactically correct script instead. If the code continues to be incorrect afterwards, we remove it and do not process it further.
To integrate the parsed scripts into existing programs, we rely on the special comments marking the beginning of sprites and scripts (cf. Section IV). Using the sprite names and block ids contained in these comments, we can align the parsed code snippets with the original code and make the required replacements. These replacements are performed on the internal program representation of LitterBox (i.e., the abstract syntax tree), which is substantially more flexible and reliable compared to working directly on the textual scratchblocks- or json-formats. Any sprites or scripts with unseen ids are added as new code to the program. When adding a new sprite, we use the default image of the Scratch cat mascot.The updated program is finally converted by LitterBox back into the Scratch-native json format and loaded into the Scratch-gui where the user can inspect it in the native code editor (upper half of Fig. 1) and continue to edit it. In case the proposed fix is not satisfactory, the user can choose to revert the change.
@@ VI Future Work
Since we introduce an initial version of our extensible LitterBox+ framework in this paper, this leaves many opportunities for future work. Our framework could be extended to other visual languages (e.g., Snap!), but suitable linters and code converters from/into text need to be developed first. For use in a classroom setting where the target solution of the program is known, additional agentic capabilities could be integrated that automatically run Whisker [9] or block-based [10] test suites so that either llm-proposed code analysis issue fixes can be checked and improved automatically via a follow-up prompt to the llm, or that fixes for failing tests can be requested directly by the user. Since the Scratch-gui is lacking any kind of code completion feature at the moment, our llm-based completion could be a valuable further extension to the framework and its users. Finally, all features proposed as part of LitterBox+ should be evaluated with children to understand how such llm-based tools are used [6, 7] to then adjust prompts and user interface, and develop new features based on the findings.
@@ VII Conclusions
Acknowledgments
We would like to thank Patric Feldmeier for his contributions to the scratchblocks parser implementation. This work is supported by DFG project FR 2955/3-3 ‘Pedagogy and Technology for Feedback in Block-Based Programming’ and FR 2955/5-1 ‘TYPES4STRINGS: Types For Strings’.